OpenStudy (anonymous):

can someone help me with this quadratic equation please? 4x(x-2)-5x(x-1)=2 If I distribute and put it in a standard form i get: -x^2-3x-2=0 and i would factor out -1 and solve x^2+3x+2 and the answer i get is {-1, -2}. The question I have is what do i do with -1 I originally factored out?

OpenStudy (anonymous):

why did you factor out -1?

OpenStudy (anonymous):

if i don't factor out -1 i cant factor my equation

OpenStudy (anonymous):

hmm i learned this backwards soo i probably can't explain how i got it im sorry

OpenStudy (anonymous):

ok thanks.....anyone else?

OpenStudy (anonymous):

share link in chat

OpenStudy (anonymous):

i have an open problem....is that the same thing?

OpenStudy (anonymous):

hold on ii'll do it

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

ignore the -1 if \(z=0\) then so does \(-z\)

OpenStudy (anonymous):

-x^2-3x-2=0...multiplying both sides by -1 gives x^2 +3x+2=0 ans..

OpenStudy (anonymous):

factoring these gives x^2 +3x+2=0 =(x+1)(x+2)=0 gives x=-2 and x=1 ans...